OpenStudy (anonymous):

A line passes through the point (0, 5) and has a slope of negative one over two. . Which is the equation of the line in slope-intercept form?

OpenStudy (anonymous):

if a line passes through x1,y1 and has a slope m then its equation is given by y-y1=m(x-x1) here x1=0 and y1=5 and slope is well stated ....